OpenStudy (anonymous):

HELP PLEASE SIMPLIFY NOT SOLVE 14j + 2j - 22k - 7j + 14k - 3k and 13g + 9g - 4 - 2g - 16g

OpenStudy (doc.brown):

What if you have 1x and another 1x, how many x's do you have?

OpenStudy (anonymous):

j(14+2-7)+k(14-3-22) -> 9j-11k Like this?

OpenStudy (anonymous):

you can just add all the j's and k's and then you have it

OpenStudy (anonymous):

14j+2j-7j=9j -22k+14k-3k=-11k

OpenStudy (doc.brown):

Do you understand why @Krokodzilla says you can add the j's?

OpenStudy (anonymous):

@doc.brown for your first comment you would have 2x and for your 2 comment yes I do and thanks @Krokodzilla

OpenStudy (doc.brown):

Yeah!

OpenStudy (anonymous):

so is the answer 11k or no? @doc.brown and @Krokodzilla

OpenStudy (doc.brown):

no

OpenStudy (doc.brown):

-22k = -k-k-k-k-k-k-k-k-k-k-k-k-k-k-k-k-k-k-k-k-k-k +14k= +k+k+k+k+k+k+k+k+k+k+k+k+k+k -3k = -k-k-k How many k's do you have?

OpenStudy (anonymous):

@doc.brown a lot but it still makes no sense Im supposed be simplifying not solving.

OpenStudy (anonymous):

If you put all the k's and j's toghether then you are indeed simplefying, you cannot solve this equation because you do not know tha value of k and j

OpenStudy (doc.brown):

"a lot"... you have -11

OpenStudy (anonymous):

So, the answer is -11?

OpenStudy (anonymous):

@doc.brown

OpenStudy (doc.brown):

The answer to how many k's you have, yes. You have -11k. \[14j + 2j - 22k - 7j + 14k - 3k\]Keep all of the same variables together\[14j + 2j - 7j - 22k + 14k - 3k\]See?\[(14j + 2j - 7j)+(- 22k + 14k - 3k)\] \[\color{red}{14j + 2j}\color{blue}{- 22k}\color{red}{- 7j}\color{blue}{+ 14k - 3k}\]\[=\color{red}{9j}\color{blue}{-11k}\]